# Non-resonant Anomaly Detection with Background Extrapolation

### v2

@@ 1 Introduction
For this reason, a new paradigm has emerged to complement traditional, model-specific approaches. The new anomaly detection protocols seek to explore data with as little bias as possible in order to be broadly sensitive to many scenarios. Such protocols have been significantly advanced by modern machine learning Kasieczka:2021xcg ; Aarrestad:2021oeb ; Karagiorgi:2021ngt ; Feickert:2021ajf  Collins:2018epr ; DAgnolo:2018cun ; Collins:2019jip ; DAgnolo:2019vbw ; Farina:2018fyg ; Heimel:2018mkt ; Roy:2019jae ; Cerri:2018anq ; Blance:2019ibf ; Hajer:2018kqm ; DeSimone:2018efk ; Mullin:2019mmh ; 1809.02977 ; Dillon:2019cqt ; Andreassen:2020nkr ; Nachman:2020lpy ; Aguilar-Saavedra:2017rzt ; Romao:2019dvs ; Romao:2020ojy ; knapp2020adversarially ; collaboration2020dijet ; 1797846 ; 1800445 ; Amram:2020ykb ; Cheng:2020dal ; Khosa:2020qrz ; Thaprasop:2020mzp ; Alexander:2020mbx ; aguilarsaavedra2020mass ; 1815227 ; pol2020anomaly ; Mikuni:2020qds ; vanBeekveld:2020txa ; Park:2020pak ; Faroughy:2020gas ; Stein:2020rou ; Chakravarti:2021svb ; Batson:2021agz ; Blance:2021gcs ; Bortolato:2021zic ; Collins:2021nxn ; Dillon:2021nxw ; Finke:2021sdf ; Shih:2021kbt ; Atkinson:2021nlt ; Kahn:2021drv ; Dorigo:2021iyy ; Caron:2021wmq ; Govorkova:2021hqu ; Kasieczka:2021tew ; Volkovich:2021txe ; Govorkova:2021utb ; Hallin:2021wme ; Ostdiek:2021bem ; Fraser:2021lxm ; Jawahar:2021vyu ; Herrero-Garcia:2021goa ; Aguilar-Saavedra:2021utu ; Tombs:2021wae ; Lester:2021aks ; Mikuni:2021nwn ; Chekanov:2021pus ; dAgnolo:2021aun ; Canelli:2021aps ; Ngairangbam:2021yma ; Bradshaw:2022qev ; Aguilar-Saavedra:2022ejy ; Buss:2022lxw ; Alvi:2022fkk ; Dillon:2022tmm ; Birman:2022xzu ; Raine:2022hht ; Letizia:2022xbe ; Fanelli:2022xwl ; Finke:2022lsu ; Verheyen:2022tov ; Dillon:2022mkq ; Caron:2022wrw ; Park:2022zov ; Kamenik:2022qxs ; Hallin:2022eoq ; Kasieczka:2022naq ; Araz:2022zxk ; Mastandrea:2022vas ; Schuhmacher:2023pro ; Roche:2023int ; Golling:2023juz ; Sengupta:2023xqy ; Mikuni:2023tok ; Golling:2023yjq ; Vaslin:2023lig ; ATLAS:2023azi ; Chekanov:2023uot ; CMSECAL:2023fvz ; Bickendorf:2023nej ; Finke:2023ltw ; Buhmann:2023acn ; Freytsis:2023cjr . One powerful anomaly detection strategy (“weak supervision”) is to isolate a region of phase space and compare data to a background-only reference. Signal model-agnostic evidence for new physics emerges when the data is not statistically consistent with the reference.
@@
The core task of weakly supervised anomaly detection strategies is to construct the reference sample of background-only events. Existing strategies follow one of two approaches: (1) the reference sample is from (background-only) simulation and (2) the reference sample is learned from sidebands Kasieczka:2021tew . The advantage of the simulation approach is that it puts few assumptions on the form of the signal. In particular, this approach can accommodate signals that are non-resonant, either due to the new physics being very massive or consisting of non-reconstructable (e.g. invisible) particles. The challenge with a simulation-based reference is that the simulation accuracy limits the anomaly detection sensitivity. In contrast, sideband methods posit that the signal, if it exists, is localized in at least one known dimension $m$. Focusing on a particular interval in $m$, a signal-sensitive region of phase space is then constructed using other features $x$, and regions in $m$ away from the posited resonance are used to estimate the background $p_{\text{background}}(x|m)$. The benefit of sideband methods is that the reference is estimated directly from data, while the challenge is that not all signals are resonant.
@@
We propose an approach to merge the advantages of the simulation-based and sideband-based cases to form a strategy that learns the reference from data without requiring a resonance. This method extends and modifies existing resonance approaches Andreassen:2020nkr ; Nachman:2020lpy ; Hallin:2021wme ; Golling:2022nkl ; Hallin:2022eoq ; 1815227 ; Sengupta:2023xqy to allow for extrapolation in $m$, as opposed to interpolation. Instead of the signal region being defined by an interval in a one-dimensional $m$, we now have a multidimensional $m$ for which the signal region is defined by thresholds [^1] A non-rectangular signal region is also possible, but makes the accounting more difficult. $m_{0}>c_{0},m_{1}>c_{1},...,m_{n}>c_{n}$ for $m,c\in\mathbb{R}^{n}$. The conditional reference $p_{\text{background}}(x|m)$ is learned (directly or indirectly) from the complement of the signal region and then extrapolated to the signal region. If the machine learning models are smooth and the background does not qualitatively change in the signal region, there is reason to believe that the extrapolation can be accurate [^2] We leave additional studies to impose minimalism in the extrapolation to future work via e.g. monotonicity Kitouni:2021fkh .. A new feature of the non-resonant case is that we need to additionally estimate $p_{\text{background}}(m)$. This is also needed in the resonant case, but it is less important (since $p(m)$ is relatively constant in the signal region) and thus approximate / simplified methods have been considered so far. We propose to estimate $p_{\text{background}}(m)$ using a likelihood-ratio method Andreassen:2020nkr and then combine this estimate with three proposals for estimating $p_{\text{background}}(m)$ based on reweighting Andreassen:2020nkr , generative models Hallin:2021wme , and template morphing Golling:2022nkl .
@@ 2.1 Locating overdensities
In the resonant case, the SR is defined by an interval around a posited new particle mass $m\in[m_{0}-\delta,m_{0}+\delta]$. The neighboring region (called sideband or SB) is the region (or a subset of the region) outside of this interval. If the signal exists and is at mass $m_{0}$, we expect that most of the signal is in the SR with little contamination in the SB. The reference background template is estimated using information from the SB. Since we expect the phase space to be smoothly varying with $m$ and since $\delta$ is expected to be small relative to regions over which large changes occur in $p_{\text{background}}(x|m)$, we expect the background to be well-estimated by interpolating from the SB (see further details in Sec. 2.2). Given samples from the background dataset [^4] There are variations on this where the density is used directly without samples Nachman:2020lpy , but the overall idea is the same., overdensities are identified with the Classification WithOut LAbels procedure (CWoLaMetodiev:2017vrx ; Collins:2018epr ; Collins:2019jip . Events in data are all assigned the label “signal” and all events from the background template are labeled “background”. An optimal classifier trained to distinguish these datasets will implicitly learn a function monotonically related to the signal / background likelihood ratio. Thresholding the classifier will then isolate an anomaly-like region of phase space.
@@ 2.2 Constructing background templates
    To circumvent these challenges, we estimate $p_{\mathrm{background}}(m)$ in the SR by reweighting context in the CR, $p_{\mathrm{simulation}}(m)$, to match data and then extrapolating this function into the SR. A binary classifier parameterized in $m$ Cranmer:2015bka ; Baldi:2016fzo is trained to distinguish simulated context from data context in the CR. The classifier output is then interpreted as a likelihood ratio $w(m)$ for estimating the background Andreassen:2020nkr . As in step (1), we assume that conditioning the network on the context will enable the learned weights to extrapolate into the SR.
@@ 2.3 Network architectures and training specifications
The Reweighting method is implemented with a binary classifier, consisting of a dense neural network with 3 layers of 100 nodes each. We use a batch size of 512, a learning rate of $10^{-3}$, and train for up to 50 epochs with an early stopping patience of 5. The Generate method is implemented with 4 layers of (Masked Affine Autoregressive Transform block of 128 hidden features, Reverse Permutation). Training is with a batch size of 216, a learning rate of $10^{-3}$, and is carried out for up to 20 epochs with an early stopping patience of 5 epochs. For the Morph method, the base density flow uses the same architecture as the Generate method. The transfer flow is similar, but layers have only $64$ hidden features, the batch size is reduced to 128, and the learning rate to $10^{-4}$.
For all anomaly detection classifier networks, we use a fully-connected architecture with 3 layers of 64 nodes each. We use a batch size of 512, a learning rate of $10^{-3}$, and train for up to 50 epochs with an early stopping patience of 5 epochs. Networks are only trained on the non-context features. In order to ensure training stability, for a given set of generated samples, we crop the weights $w(m)$ by dropping all weights that are greater than 3 standard deviations above the mean. This has the effect of removing a very small number (usually on the order of a few tens, or $\sim 0.1\%$) of samples whose weights severely bias the resulting $p_{\rm background}(x,m)$ distributions.
@@ 3.2.1 Closure tests
As a first test of extrapolation, we check that we are able to generate realistic background samples in both the CR and SR under zero-signal injection. In Fig. 4, we show the five feature distributions as constructed by the Reweight, Generate, and Morph methods in relation to truth, the zero-signal data. Feature reconstruction is excellent in the CR, and the ratios of the marginals to the truth are consistent with unity across virtually all of the domains of the observables. Reconstruction is good overall in the SR, although the reconstructed distributions (particularly for the Morph method) deviate visibly.
@@
In Fig. 4(a), we show a more rigorous test by training a binary classifier to discriminate between the constructed samples and background-only data. We show the spread of ROC AUCs (receiver operating characteristic area-under-curves) across $10$ binary classifier runs; for a classifier trained on two identical datasets, the ROC AUC is expected to be 0.5 for a perfect classifier, in the limit of infinite statistics and training time. Again, we see that there is excellent reconstruction in the CR and slightly worse reconstruction in the SR. In Fig. 4(b), we show the classifier rejection (1 / false positive rate) curves for the discrimination task in the SR. For all of the methods, the rejection found in the zero-signal case is compatible (to within errorbars) with that of a random classifier.
@@ 3.2.2 Signal detection tests
In this section, we carry out an anomaly detection test for the dark QCD physics model. The test is done by generating a signal-background discriminator using a weekly-supervised binary classifier following the CWoLa procedure, as explained in Sec. 2. The classifier is trained to discriminate the background reference against data that consists of SR background with a known amount of signal injection. For the simulation, we scan over seven signal injections of signal ($S$) over background ($B$) of $S/B\in\{0,3.1,6.5,9.3,12,15,18\}\times 10^{-3}$ [^8] Note that with our chosen context region bounds, there is a small amount of signal contamination in the CR. For the largest signal injection, the contamination is 4.3×$\times 10^{-3}$%, from 384 events., corresponding to significances $S/\sqrt{B}\in\{0,0.84,1.75,2.52,3.32,4.07,4.99$}, respectively. In this section, we provide a number of metrics derived from this binary classifier.
@@ Appendix B Modifying the simulation
We introduce a hand-tuned modification to the simulation, both to the features and context, defined in Eq. 2:

### v3

@@ 1 Introduction
For this reason, a new paradigm has emerged to complement traditional, model-specific approaches. The new anomaly detection protocols seek to explore data with as little bias as possible in order to be broadly sensitive to many scenarios. Such protocols have been significantly advanced by modern machine learning (see Kasieczka:2021xcg ; Aarrestad:2021oeb ; Karagiorgi:2021ngt ; Feickert:2021ajf for a selection of machine learning reviews and anomaly detection challenges, and Collins:2018epr ; DAgnolo:2018cun ; Collins:2019jip ; DAgnolo:2019vbw ; Farina:2018fyg ; Heimel:2018mkt ; Roy:2019jae ; Cerri:2018anq ; Blance:2019ibf ; Hajer:2018kqm ; DeSimone:2018efk ; Mullin:2019mmh ; 1809.02977 ; Dillon:2019cqt ; Andreassen:2020nkr ; Nachman:2020lpy ; Aguilar-Saavedra:2017rzt ; Romao:2019dvs ; Romao:2020ojy ; knapp2020adversarially ; collaboration2020dijet ; 1797846 ; 1800445 ; Amram:2020ykb ; Cheng:2020dal ; Khosa:2020qrz ; Thaprasop:2020mzp ; Alexander:2020mbx ; aguilarsaavedra2020mass ; 1815227 ; pol2020anomaly ; Mikuni:2020qds ; vanBeekveld:2020txa ; Park:2020pak ; Faroughy:2020gas ; Stein:2020rou ; Chakravarti:2021svb ; Batson:2021agz ; Blance:2021gcs ; Bortolato:2021zic ; Collins:2021nxn ; Dillon:2021nxw ; Finke:2021sdf ; Shih:2021kbt ; Atkinson:2021nlt ; Kahn:2021drv ; Dorigo:2021iyy ; Caron:2021wmq ; Govorkova:2021hqu ; Kasieczka:2021tew ; Volkovich:2021txe ; Govorkova:2021utb ; Hallin:2021wme ; Ostdiek:2021bem ; Fraser:2021lxm ; Jawahar:2021vyu ; Herrero-Garcia:2021goa ; Aguilar-Saavedra:2021utu ; Tombs:2021wae ; Lester:2021aks ; Mikuni:2021nwn ; Chekanov:2021pus ; dAgnolo:2021aun ; Canelli:2021aps ; Ngairangbam:2021yma ; Bradshaw:2022qev ; Aguilar-Saavedra:2022ejy ; Buss:2022lxw ; Alvi:2022fkk ; Dillon:2022tmm ; Birman:2022xzu ; Raine:2022hht ; Letizia:2022xbe ; Fanelli:2022xwl ; Finke:2022lsu ; Verheyen:2022tov ; Dillon:2022mkq ; Caron:2022wrw ; Park:2022zov ; Kamenik:2022qxs ; Hallin:2022eoq ; Kasieczka:2022naq ; Araz:2022zxk ; Mastandrea:2022vas ; Schuhmacher:2023pro ; Roche:2023int ; Golling:2023juz ; Sengupta:2023xqy ; Mikuni:2023tok ; Golling:2023yjq ; Vaslin:2023lig ; ATLAS:2023azi ; Chekanov:2023uot ; CMSECAL:2023fvz ; Bickendorf:2023nej ; Finke:2023ltw ; Buhmann:2023acn ; Freytsis:2023cjr for specific techniques). One powerful anomaly detection strategy (“weak supervision”) is to isolate a region of phase space and compare data to a background-only reference. Signal model-agnostic evidence for new physics emerges when the data is not statistically consistent with the reference.
@@
The core task of weakly supervised anomaly detection strategies is to construct the reference sample of background-only events. Existing strategies follow one of two approaches: (1) the reference sample is from (background-only) simulation and (2) the reference sample is learned from sidebands. The advantage of the simulation approach is that it puts few assumptions on the form of the signal. In particular, this approach can accommodate signals that are non-resonant, either due to the new physics being very massive or consisting of non-reconstructable (e.g. invisible) particles. The challenge with a simulation-based reference is that the simulation accuracy limits the anomaly detection sensitivity. In contrast, sideband methods posit that the signal, if it exists, is localized in at least one known dimension $m$. Focusing on a particular interval in $m$, a signal-sensitive region of phase space is then constructed using other features $x$, and regions in $m$ away from the posited resonance are used to estimate the background $p_{\text{background}}(x|m)$. The benefit of sideband methods is that the reference is estimated directly from data, while the challenge is that not all signals are resonant.
@@
We propose an approach to merge the advantages of the simulation-based and sideband-based cases to form a strategy that learns the reference from data without requiring a resonance. This method extends and modifies existing resonance approaches Andreassen:2020nkr ; Nachman:2020lpy ; Hallin:2021wme ; Golling:2022nkl ; Hallin:2022eoq ; 1815227 ; Sengupta:2023xqy to allow for extrapolation in $m$, as opposed to interpolation. Instead of the signal region being defined by an interval in a one-dimensional $m$, we now have a multidimensional $m$ for which the signal region is defined by thresholds [^1] A non-rectangular signal region is also possible, but makes the accounting more difficult. $m_{0}>c_{0},m_{1}>c_{1},...,m_{n}>c_{n}$ for $m,c\in\mathbb{R}^{n}$. The conditional reference $p_{\text{background}}(x|m)$ is learned (directly or indirectly) from the complement of the signal region and then extrapolated to the signal region. If the machine learning models are smooth and the background remains qualitatively the same in the signal region (e.g. the functional form of the background distribution, particularly the dependence on the context variables, is continuous across the CR-SR boundary), there is reason to believe that the extrapolation can be accurate [^2] We leave additional studies to impose minimalism in the extrapolation to future work via e.g. monotonicity Kitouni:2021fkh .. A new feature of the non-resonant case is that we need to additionally estimate $p_{\text{background}}(m)$. This is also needed in the resonant case, but it is less important (since $p(m)$ is relatively constant in the signal region) and thus approximate or simplified methods have been considered so far. We propose to estimate $p_{\text{background}}(x|m)$ using a likelihood-ratio method Andreassen:2020nkr and then combine this estimate with three proposals for estimating $p_{\text{background}}(m)$ based on reweighting Andreassen:2020nkr , generative models Hallin:2021wme , and template morphing Golling:2022nkl .
@@
Non-resonant signals can arise from many new physics models. For example, models of a strongly interacting dark sector can produce jets which contain both stable and unstable dark hadrons, predicting signatures of jets containing significant missing energy. Non-resonant signals are also characteristic of effective field theories, which can model new physics with masses beyond those accessible through on-shell production. ML-based anomaly detection can offer additional sensitivity to explore the vast signature space of non-resonant signals.
@@ 2.1 Locating overdensities
In the resonant case, the SR is defined by an interval around a posited new particle mass $m\in[m_{0}-\delta,m_{0}+\delta]$. The neighboring region (called sideband or SB) is the region (or a subset of the region) outside of this interval. If the signal exists and is at mass $m_{0}$, we expect that most of the signal is in the SR with little contamination in the SB. The reference background template is estimated using information from the SB. Since we expect the phase space to be smoothly varying with $m$ and since $\delta$ is expected to be small relative to regions over which large changes occur in $p_{\text{background}}(x|m)$, we expect the background to be well-estimated by interpolating from the SB (see further details in Sec. 2.2). Given samples from the background dataset [^4] There are variations on this where the density is used directly without samples Nachman:2020lpy , but the overall idea is the same., overdensities are identified with the Classification WithOut LAbels procedure (CWoLaMetodiev:2017vrx ; Collins:2018epr ; Collins:2019jip . Events in data are all assigned the label “signal” and all events from the background template are labeled “background”. An optimal classifier trained to distinguish these datasets will implicitly learn a function monotonically related to the signal-over-background likelihood ratio. Thresholding the classifier will then isolate an anomaly-like region of phase space.
@@ 2.2 Constructing background templates
    To circumvent these challenges, we estimate $p_{\mathrm{background}}(m)$ in the SR by reweighting context in the CR, $p_{\mathrm{simulation}}(m)$, to match data and then extrapolating this function into the SR. A binary classifier parameterized in $m$ Cranmer:2015bka ; Baldi:2016fzo is trained to distinguish simulated context from data context in the CR. The classifier output is then interpreted as a likelihood ratio $w(m)$ for estimating the background Andreassen:2020nkr . As in step 1, we assume that conditioning the network on the context will enable the learned weights to extrapolate into the SR.
@@ 2.3 Network architectures and training specifications
The Reweighting method is implemented with a binary classifier, consisting of a dense neural network with 3 layers of 100 nodes each. We use a batch size of 512, a learning rate of $10^{-3}$, and train for up to 50 epochs with an early stopping patience of 5. The classifier that generates the context weights $w(m)$ is exactly the same as that for the Reweighting method. The Generate method is implemented with 4 layers of (Masked Affine Autoregressive Transform block of 128 hidden features, Reverse Permutation), where the transform block is built from a Masked Autoencoder for Distribution Estimation (MADE) architecture. Training is with a batch size of 216, a learning rate of $10^{-3}$, and is carried out for up to 20 epochs with an early stopping patience of 5 epochs. For the Morph method, the base density flow uses the same architecture as the Generate method. The transfer flow is similar, but layers have only $64$ hidden features, the batch size is reduced to 128, and the learning rate to $10^{-4}$.
For the anomaly detection classifier networks, as well as the networks used for the closure tests (see Sec. 3.2.1), we use a fully-connected architecture with 3 layers of 64 nodes each. We use a batch size of 512, a learning rate of $10^{-3}$, and train for up to 50 epochs with an early stopping patience of 5 epochs. Networks are only trained on the non-context features. In order to ensure training stability, for a given set of generated samples, we crop the weights $w(m)$ by dropping all weights that are greater than 3 standard deviations above the mean. This has the effect of removing a very small number (usually on the order of a few tens, or $\sim 0.1\%$) of samples whose weights severely bias the resulting $p_{\rm background}(x,m)$ distributions.
@@ 3.2.1 Closure tests
As a first test of extrapolation, we check that we are able to generate realistic background samples in both the CR and SR under zero-signal injection. In Fig. 4, we show the five feature distributions as constructed by the Ideal, Reweight, Generate, and Morph methods in relation to truth, the zero-signal data. Feature reconstruction is excellent in the CR, and the ratios of the marginals to the truth are consistent with unity across virtually all of the domains of the observables. Reconstruction is good overall in the SR, although the reconstructed distributions (particularly for the Morph method) deviate visibly. We calculated the one-dimensional Wasserstein distances as a quantitative measurement of the closure test, characterizing how well the observables are reconstructed in the signal region and control region by Reweight, Generate and Morph methods. The average Wasserstein distances of Ideal, Reweight, Generate, and Morph across five feature variables are $0.0048$, $0.0073$, $0.0237$, $0.0116$ in CR, and $0.0058$, $0.0130$, $0.0293$, $0.0202$ in SR, respectively. Note that in order to avoid premature unblinding, we withhold a set of 10k simulation and 10k data events in the CR from network training and use this set to evaluate CR closure.
@@
[…]
@@ 3.2.2 Signal detection tests
In this section, we carry out an anomaly detection test for the dark QCD physics model. The test is done by generating a signal-background discriminator using a weekly-supervised binary classifier following the CWoLa procedure, as explained in Sec. 2. The classifier is trained to discriminate the background reference against data that consists of SR background with a known amount of signal injection. For the simulation, we scan over seven signal injections of signal ($S$) over background ($B$) of $S/B\in\{0,3.1,6.5,9.3,12,15,18\}\times 10^{-3}$, corresponding to significances $S/\sqrt{B}\in\{0,0.84,1.75,2.52,3.32,4.07,4.99$}, respectively [^8] Note that with our chosen context region bounds, there is a small amount of signal contamination in the CR. For the largest signal injection, the contamination is 4.3×$\times 10^{-3}$%, from 384 events.. In this section, we provide a number of metrics derived from this binary classifier.
@@ Appendix B Modifying the simulation
We introduce a hand-tuned modification to the simulation, both to the features and context, defined in Eq. LABEL:eq:sim_mod: